OpenStudy (ksaimouli):

what is area of circle whose radius is 4cm

OpenStudy (anonymous):

do you know the formula for the area of a circle with radius R?

OpenStudy (ksaimouli):

ya ans

OpenStudy (ksaimouli):

answ

OpenStudy (anonymous):

= pi*r^2 =22/7*(4)^2

OpenStudy (anonymous):

=(22/7)*16

OpenStudy (ksaimouli):

i know just answer

OpenStudy (anonymous):

50.29